An 18.3 kg potted plant is put on a 24.4 kg wooden table. The table and the floor do not budge. a. What is the normal force of the table on the potted plant?

b. What is the normal force of the floor on all of the table legs in total?


1
Expert's answer
2021-01-28T15:45:36-0500

a)


"N_1=(m_1)g\\N_1=(18.3)9.8=179\ N"

b)


"N_2=(m_1+m_2)g\\N_2=(18.3+24.4)9.8=418\ N"